OpenStudy (anonymous):

Find the midpoint between two points on a coordinate plane whose coordinates are (7, -8) and (-3, -4). please help medal and fanning in order;)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Midpoint is just the average of the x coordinates and the average of the y coordinates.

OpenStudy (anonymous):

could you show me the steps of doing this problem?

OpenStudy (whpalmer4):

Average of 7 and -3 is? Average of -8 and -4 is? Those are the x,y coordinates.

OpenStudy (anonymous):

x=x(sub1) + x(sub2)/2 and y=y(sub1) +y(sub2)/2

OpenStudy (whpalmer4):

My 10-year-old wants you to know that he wants me to help him on his science fair project and that the answer is (2,-6) because 7 + -3 = 4 and 4/2 = 2, and -8 + -4 = -12 and -12/2 = -6

OpenStudy (whpalmer4):

Being 10, he's too young for an openstudy account, so I have to be his amanuensis :-)